OpenStudy (anonymous):

Find the flux of F = x^3i  + y^3 j  + z^3 k  out of the closed surface z = squroot (x^2 + y^2) a cone 0 ≤ z ≤ 4 oriented outward.

OpenStudy (irishboy123):

divergence theorem?? personal pref would be to set it up in cartesian as \(\iiint \, dx \, dy \, dz \quad div (\vec F) = \iiint dx\, dy \, dz \quad 3(x^2 + y^2 + z^2) \) then switch to cylindrical: \(3\int_{0}^{2\pi}\, \int_{0}^{4}\, \int_{0}^{z}\, dr \, dz \, d\theta \quad (r^2 + z^2)r \) or \(3\int_{0}^{2\pi}\, \int_{0}^{4}\, \int_{r}^{4}\, dz \, dr \, d\theta \quad (r^2 + z^2)r \).